The measure of the smallest angle of a triangle is
​one-fifth
the measure of the largest angle. The measure of the second angle is
95 degrees°
less than the measure of the largest angle. Find the measure of the angles of the triangle.

Answer by stanbon(75887)

The measure of the smallest angle of a triangle is
​one-fifth the measure of the largest angle.
The measure of the second angle is 95 degrees° less than the measure of the largest angle.
Find the measure of the angles of the triangle.
------
largest angle:: x
smallest angle:: x/5
2nd angle:: x-95
-------------------
Equation:
x + (x/5) + x-95 = 180
(11/5)x = 275
x = 25*5 = 125 degrees
x/5 = 25 degrees
x-95 = 30 degrees
--------------
Cheers
Stan H.
---------